# Finite simple groups acting with fixity $4$ and their occurrence as groups of automorphisms of Riemann surfaces

### v1

@@ 4 Proof of the main result
For $\textrm{M}_{11}$, we have one list, and a suitable generating tuple is
(((1,4,11,10,7,5,8,6,9,3,2)),((3,7,4,5)(6,11,9,8)),(((1,4,11,10,7,5,8,6,9,3,2)),((3,7,4,5)(6,11,9,8)),( ( ( 1 , 4 , 11 , 10 , 7 , 5 , 8 , 6 , 9 , 3 , 2 ) ) , ( ( 3 , 7 , 4 , 5 ) ( 6 , 11 , 9 , 8 ) ) ,
((2,5,4,7,6)(3,11,9,8,10)))((2,5,4,7,6)(3,11,9,8,10)))( ( 2 , 5 , 4 , 7 , 6 ) ( 3 , 11 , 9 , 8 , 10 ) ) ) for $[\textrm{M}_{11},g,1\mid[5,1]]$.
There are two lists for $\textrm{M}_{22}$:
The first generating tuple we found is
(((1,11,7,19,21,20,9,15,8,3,2)(4,10,6,22,16,14,12,17,18,5,13)),((1,7,11,13,22,6,9)(2,18,21,3,14,17,20)(5,8,10,12,19,16,15)),((3,14,6,15,22)(4,16,7,5,11)(8,20,21,9,17)(10,13,19,18,12)))111719212091583241062216141217185131711132269218213141720581012191615314615224167511820219171013191812(((1,11,7,19,21,20,9,15,8,3,2)(4,10,6,22,16,14,12,17,18,5,13)),\\ ((1,7,11,13,22,6,9)(2,18,21,3,14,17,20)(5,8,10,12,19,16,15)),\\ ((3,14,6,15,22)(4,16,7,5,11)(8,20,21,9,17)(10,13,19,18,12)))( ( ( 1 , 11 , 7 , 19 , 21 , 20 , 9 , 15 , 8 , 3 , 2 ) ( 4 , 10 , 6 , 22 , 16 , 14 , 12 , 17 , 18 , 5 , 13 ) ) , ( ( 1 , 7 , 11 , 13 , 22 , 6 , 9 ) ( 2 , 18 , 21 , 3 , 14 , 17 , 20 ) ( 5 , 8 , 10 , 12 , 19 , 16 , 15 ) ) , ( ( 3 , 14 , 6 , 15 , 22 ) ( 4 , 16 , 7 , 5 , 11 ) ( 8 , 20 , 21 , 9 , 17 ) ( 10 , 13 , 19 , 18 , 12 ) ) )
for $[\textrm{M}_{22},g,1\mid[5,1]]$. The second is
(((1,11,7,19,21,20,9,15,8,3,2)(4,10,6,22,16,14,12,17,18,5,13)),((1,7,11,13,22,6,9)(2,18,21,3,14,17,20)(5,8,10,12,19,16,15)),((2,17,9,13,12)(4,8,18,20,16)(5,19,6,10,14)(7,11,22,15,21),(2,10,15,3,14,13,17)(4,7,9,8,16,21,22)(5,6,18,20,12,19,11)))1117192120915832410622161412171851317111322692182131417205810121916152179131248182016519610147112215212101531413174798162122561820121911(((1,11,7,19,21,20,9,15,8,3,2)(4,10,6,22,16,14,12,17,18,5,13)),\\ ((1,7,11,13,22,6,9)(2,18,21,3,14,17,20)(5,8,10,12,19,16,15)),\\ ((2,17,9,13,12)(4,8,18,20,16)(5,19,6,10,14)(7,11,22,15,21),\\ (2,10,15,3,14,13,17)(4,7,9,8,16,21,22)(5,6,18,20,12,19,11)))( ( ( 1 , 11 , 7 , 19 , 21 , 20 , 9 , 15 , 8 , 3 , 2 ) ( 4 , 10 , 6 , 22 , 16 , 14 , 12 , 17 , 18 , 5 , 13 ) ) , ( ( 1 , 7 , 11 , 13 , 22 , 6 , 9 ) ( 2 , 18 , 21 , 3 , 14 , 17 , 20 ) ( 5 , 8 , 10 , 12 , 19 , 16 , 15 ) ) , ( ( 2 , 17 , 9 , 13 , 12 ) ( 4 , 8 , 18 , 20 , 16 ) ( 5 , 19 , 6 , 10 , 14 ) ( 7 , 11 , 22 , 15 , 21 ) , ( 2 , 10 , 15 , 3 , 14 , 13 , 17 ) ( 4 , 7 , 9 , 8 , 16 , 21 , 22 ) ( 5 , 6 , 18 , 20 , 12 , 19 , 11 ) ) )
for $[\textrm{M}_{22},g,1\mid[5,1],[7,1]]$.
We calculated a generating tuple for $\textrm{J}_{1}$ with the GAP code in the next section, but again we do not display it here because of its size.
@@ 5 The GAP code
local ct,cc,list,pos,ro,tuple,orb,orb1,h,i,len;
@@
local ct,cc,list,pos,ro,tuple,orb,orb1,h,i,len;
@@
To compute a generating tuple for the group $\textrm{Alt}_{7}$ and for the Hurwitz datum $[\textrm{Alt}_{7},g,g_{0}\mid[5,1]]$, we run GenTuple(AlternatingGroup(7),1,[5]);
For the list and $[\textrm{Alt}_{7},g,g_{0}\mid[5,1],[7,1]]$ we use a different approach, because GenTuple is very slow. This is how we obtained the generating tuple in the proof of Theorem 1.1:
@@
For $\textrm{PSL}_{2}(7)$ we have six different lists to check. The ones that can be treated with the function GenTuple are $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1]]$, $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[7,1]]$ and $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$. Bearing in mind the smallest value for $g_{0}$, we use, after setting $G$ to $\textrm{PSL}_{2}(7)$, the commands
@@
A generating tuple for the list $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$ can be used to construct a tuple for the list $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2],[7,1]]$ as follows:
@@
Now we look at two cases that we treat differently. The first list in the table is $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1]]$, and we check it for the smallest possible value of $g_{0}$ as follows:
@@
For the list $[\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2]]$, we also do something different because GenTuple is very slow here:
@@
For $\textrm{PSL}_{2}(8)$ we have even more cases to consider. First we look at the lists that can be treated with GenTuple, again with the smallest cogenus, namely $[\textrm{PSL}_{2}(8),g,1\mid[2,1],[7,1]]$, $[\textrm{PSL}_{2}(8),g,1\mid[2,1],[9,1]]$,
$[\textrm{PSL}_{2}(8),g,0\mid[2,1],[9,2]]$ and $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,1],[9,1]]$.
@@
The list $[\textrm{PSL}_{2}(8),g,2\mid[2,1]]$ is treated differently:
@@
Finally, we construct a generating tuple for the list $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2]]$. We must treat this differently from just running the function GenTuple, because in the definition of the variable ro in GenTuple we pick a random number. But for the list $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2]]$, things go wrong if ro[2] and ro[3] are accidentally picked to be the same number within GenTuple. Therefore, we add an additional step to make sure that ro[2] and ro[3] are different.
@@
cc := ConjugacyClasses(G);;
cco := List(cc, i-> Order(Representative(i)));;
@@
For the group $\textrm{PSU}_{4}(3)$ we have two lists in our table. Since GenTuple takes a long time here, we construct our generating tuples differently. First we look at the list $[\textrm{PSU}_{4}(3),g,1\mid[5,1]]$, we calculate a candidate and check whether it satisfies all conditions.
@@
For the second list, $[\textrm{PSU}_{4}(3),g,1\mid[5,1],[7,1]]$, we start similarly and then change the generators slightly:
@@
Finally, we have three sporadic groups to consider, with four lists in total. GenTuple is often too slow here. For $\textrm{M}_{11}$ and the list $[\textrm{M}_{11},g,1\mid[5,1]]$, we use the following code:
@@
The next lists in the table are for the group $\textrm{M}_{22}$. First we look at $[\textrm{M}_{22},g,1\mid[5,1]]$:
@@
Next we consider $[\textrm{M}_{22},g,1\mid[5,1],[7,1]]$, again using some of the information from before.
@@
Finally, we have the code for the list $[\textrm{J}_{1},g,g_{0}\mid[15,1]]$. Here we use an additional GAP package, called TomLib. The TomLib package is only used for the construction of the group $\textrm{J}_{1}$ as there is no predefined function in GAP to construct this group. After we generate $\textrm{J}_{1}$ we proceed in a similar way as before.

### v2

@@ 4 Proof of the main result
For $\textrm{M}_{11}$, we have one list, namely $\textrm{M}_{11},g,1\mid[5,1]]$, and GAP give us a suitable generating tuple:
(((1,4,11,10,7,5,8,6,9,3,2)),((3,7,4,5)(6,11,9,8)),(((1,4,11,10,7,5,8,6,9,3,2)),((3,7,4,5)(6,11,9,8)),
$((2,5,4,7,6)(3,11,9,8,10)))$.
There are two lists for $\textrm{M}_{22}$ and one list for $\textrm{J}_{1}$, and again we do not display the generating tuples that we found with GAP. The group $\textrm{J}_{1}$ is the only one that we treat differently from the others, by using the TomLib package ([16]) and constructing the group from there. This concludes the proof of Theorem 1.1.
@@ 5 The GAP code
local ct,$c$,list,pos,ro,tuple,orb,orb1,h,i,len;
@@
local ct,$c$,list,pos,ro,tuple,orb,orb1,h,i,len;
@@
To compute a generating tuple for the group $\textrm{Alt}_{7}$ and for the Hurwitz datum $\textrm{Alt}_{7},g,g_{0}\mid[5,1]]$, we run GenTuple(AlternatingGroup(7),1,[5]);
For the list and $\textrm{Alt}_{7},g,g_{0}\mid[5,1],[7,1]]$ we use a different approach, because GenTuple is very slow. This is how we obtained the generating tuple in the proof of Theorem 1.1:
@@
For $\textrm{PSL}_{2}(7)$ we have six different lists to check. The ones that can be treated with the function GenTuple are $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1]]$, $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[7,1]]$ and $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$. Bearing in mind the smallest value for $g_{0}$, we use, after setting $G$ to $\textrm{PSL}_{2}(7)$, the commands
@@
A generating tuple for the list $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,1],[7,1]]$ can be used to construct a tuple for the list $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2],[7,1]]$ as follows:
@@
Now we look at two cases that we treat differently. The first list in the table is $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1]]$, and we check it for the smallest possible value of $g_{0}$ as follows:
@@
For the list $\textrm{PSL}_{2}(7),g,g_{0}\mid[2,1],[3,2]]$, we also do something different because GenTuple is very slow here:
@@
For $\textrm{PSL}_{2}(8)$ we have even more cases to consider. First we look at the lists that can be treated with GenTuple, again with the smallest cogenus, namely $\textrm{PSL}_{2}(8),g,1\mid[2,1],[7,1]]$, $\textrm{PSL}_{2}(8),g,1\mid[2,1],[9,1]]$,
$\textrm{PSL}_{2}(8),g,0\mid[2,1],[9,2]]$ and $\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,1],[9,1]]$.
@@
The list $\textrm{PSL}_{2}(8),g,2\mid[2,1]]$ is treated differently:
@@
Finally, we construct a generating tuple for the list $\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2]]$. We must treat this differently from just running the function GenTuple, because in the definition of the variable ro in GenTuple we pick a random number. But for the list $\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2]]$, things go wrong if ro[2] and ro[3] are accidentally picked to be the same number within GenTuple. Therefore, we add an additional step to make sure that ro[2] and ro[3] are different.
@@
$c$ := ConjugacyClasses(G);;
cco := List($c$, i-> Order(Representative(i)));;
@@
For the group $\textrm{PSU}_{4}(3)$ we have two lists in our table. Since GenTuple takes a long time here, we construct our generating tuples differently. First we look at the list $\textrm{PSU}_{4}(3),g,1\mid[5,1]]$, we calculate a candidate and check whether it satisfies all conditions.
@@
For the second list, $\textrm{PSU}_{4}(3),g,1\mid[5,1],[7,1]]$, we start similarly and then change the generators slightly:
@@
Finally, we have three sporadic groups to consider, with four lists in total. GenTuple is often too slow here. For $\textrm{M}_{11}$ and the list $\textrm{M}_{11},g,1\mid[5,1]]$, we use the following code:
@@
The next lists in the table are for the group $\textrm{M}_{22}$. First we look at $\textrm{M}_{22},g,1\mid[5,1]]$:
@@
Next we consider $\textrm{M}_{22},g,1\mid[5,1],[7,1]]$, again using some of the information from before.
@@
Finally, we have the code for the list $\textrm{J}_{1},g,g_{0}\mid[15,1]]$. Here we use an additional GAP package, called TomLib. The TomLib package is only used for the construction of the group $\textrm{J}_{1}$ as there is no predefined function in GAP to construct this group. After we generate $\textrm{J}_{1}$ we proceed in a similar way as before.